OpenStudy (anonymous):

The expression (2x + 1)(3x + 4) simplifies to which of the following? A. 2x(3x + 4) B. 2x(3x + 4) + 3(3x + 4) C. 2x(3x + 4) + (3x + 4) D. 2x(2x + 3) + 3(3x + 4)

OpenStudy (dan815):

Expand the brackets like so (a+b)(c)= ac + bc , now in your case c=3x+4 and a+b = 2x+1 (2x+1)*(3x+4)= 2x(3x+4)+1*(3x+4), can you find one that is very similar to this in your answers

OpenStudy (anonymous):

C?

OpenStudy (dan815):

yep